OpenStudy (anonymous):

how do u do -z + 3 = 7

OpenStudy (amistre64):

Are you looking for help to understand HOW to do these types of problems? or just wanting someone else to give you the answers?

OpenStudy (anonymous):

im looking for help to understand how to do the problems

OpenStudy (amistre64):

Then lets start with what we know about an equation. there is a lefthand side (LHS) and a righthand side (RHS) seperated by a sign that indicates that the LHS is equal in value to the RHS. Think of these sides as the two sides of a scale. If they are equal to each other, than the scales ar balanced. Whatever you add or subtract from one side has to be done to the other side for them to remain balanced. easy example: 10 = 10. If we alter the LHS by adding 5 to it; we have to add 5 to the RHS in order to keep things balanced. That is the key to solving your equation here. -z +3 = 7 we need to find out what "z" is equal to. lets subtract (3) from each side: -z +3 -3 = 7 -3 -z = 4 We know that a (-) times a (-) gives us a (+); and also that anything times (1) equals itself. So lets multiply each side by (-1) and see what happens. -z(-1) = 4(-1) z=-4 Lets double check our answer by putting the value of (-4) in place of (z) in the original equation: -(-4) + 3 = 7 Remember that a (-) times a (-) equals a (+) so: 4 + 3 = 7 7 = 7 is a true statement so we know that the answer is: z= (-4)